OpenStudy (anonymous):

7t^2-28t/2t^2-5t-12 times 6t^2-t-15/49t^3

OpenStudy (anonymous):

\[7t^2-28t/2t^2-5t-12 \times 6t^2-t-15/49t^3\]

OpenStudy (anonymous):

first fraction 7t(t-4)/(2t + 3)(t - 4 ) 7t/2t+3

OpenStudy (anonymous):

second fracton (2t + 3)(3t- 5 )/49t^3

OpenStudy (anonymous):

\[\frac{7t}{2t+3} \frac{ (2t + 3)(3t- 5 )}{49t^3}\]

OpenStudy (anonymous):

\[7t \frac{ (3t- 5 )}{49t^3}\]

OpenStudy (anonymous):

Ok, thank you for clarifying, and that simplifies to..?

OpenStudy (anonymous):

\[ \frac{ (3t- 5 )}{7t^2}\]

OpenStudy (anonymous):

Ok, thanks.